# Dose-aware Diffusion Model for 3D Low-count PET: Multi-institutional Validation with Reader Study and Real Low-dose Data

### v2

@@ IV Discussion and Conclusion
Compared to previous diffusion models which are time-consuming to reconstruct the entire 3D volumes, once the DDPET-3D is fully-trained, the denoising process is efficient and can be finished within 15 min using one NVIDIA A40 GPU.
Compared to most of the previously published PET denoising methods [11, 12, 6, 7, 8, 9, 10, 13, 14, 15, 17], which learn the mapping from images collected at a specific dose-level to the normal-dose counterparts, the proposed DDPET-3D can be viewed as a major extension that learns to achieve dose-aware denoising by training the network using dataset with a wide range of input count-levels and associating injected dose within the network. Since the noise-level in PET images can be affected by various factors in clinical settings, the ability to generalize to different noise-levels would be beneficial to implement the algorithm for routine clinical use.
Confirmed with reader studies by professional nuclear physicians, the proposed DDPET-3D can provide a similar or better image quality compared with normal-dose images in terms of overall subjective visual impression. DDPET-3D effectively suppressed background noise and false hot spots in certain patients, making true lesions/hot spots easier to identify.
Because of the above-mentioned benefits, DDPET-3D can be easily integrated with current clinical reconstruction workflow to denoise PET images.
@@
In conclusion, the proposed DDPET-3D provides better or similar image quality in different low-dose settings compared to images acquired in a normal-dose setting as measured by an expert visual preference study. The experimental results showed that DDPET-3D achieved superior performance compared with previous methods. While this study focuses on PET image denoising, we believe that the proposed method could be easily extended to other 3D reconstruction tasks for different imaging modalities.
@@ Acknowledgment
Ramsey D. Badawa, Benjamin A. Spencer, and Reimund Bayerlein declared that UC Davis has a revenue sharing agreement with United Imaging Healthcare.

### v3

@@ V Discussion and Conclusion
Compared to previous diffusion models which are time-consuming to reconstruct the entire 3D volumes, once the DDPET-3D is fully trained, the denoising process is efficient and can be finished within 15 min using one NVIDIA A40 GPU.
Compared to most of the previously published PET denoising methods [11, 12, 6, 7, 8, 9, 10, 13, 14, 15, 17], which learn the mapping from images collected at a specific dose level to the normal dose counterparts, the proposed DDPET-3D can be viewed as a major extension that learns to achieve dose-aware denoising by training the network using dataset with a wide range of input count levels and associating injected dose within the network. Since the noise level in PET images can be affected by various factors in clinical settings, the ability to generalize to different noise levels would be beneficial to implement the algorithm for routine clinical use.
Confirmed with reader studies by professional nuclear physicians, the proposed DDPET-3D can provide a similar or better image quality compared to normal dose images in terms of overall subjective visual impression. DDPET-3D effectively suppressed background noise and false hot spots in certain patients, making true lesions/hot spots easier to identify.
Due to the benefits mentioned above, DDPET-3D can be easily integrated with current clinical reconstruction workflow to denoise PET images.
@@
In conclusion, the proposed DDPET-3D provides better or similar visual image quality in different low-dose settings compared to images acquired in a normal-dose setting as measured by an expert visual preference study. The experimental results showed that DDPET-3D achieved superior performance compared to previous methods. Although this study focuses on PET image denoising, we believe that the proposed technical innovations could easily be extended to other medical imaging problems, as well as to video processing and general computer vision problems involving volumetric or sequential data. In particular, the importance of maintaining slice consistency in 3D medical imaging problems is analogous to maintaining temporal consistency in video problems.
@@ Acknowledgment
Ramsey D. Badawi, Benjamin A. Spencer, and Reimund Bayerlein declared that UC Davis has a revenue sharing agreement with United Imaging Healthcare.